# arXiv:2602.23189v2[math.AP] 02 Mar 2026

### v2

@@ 1.3 Main theorem
Assumptions (1.9) corresponds to the standard notion of well-prepared initial data for the singular limit, as in [36]. They ensure that the initial deviations from the incompressible limit are sufficiently small to control the energy estimates uniformly in $\varepsilon$.
Remark 1.8.
The assumption (1.6) is preserved along the time by conservation of mass:
$\int_{\Omega}R_{\pm}^{\varepsilon}(t,x)\,dx=\int_{\Omega}R_{\pm,0}^{\varepsilon}(x)\,dx,\qquad\int_{\Omega}R_{\pm}(t,x)\,dx=\int_{\Omega}R_{\pm,0}(x)\,dx,\qquad\text{for all }t\in(0,T).$
The analysis relies on the relative entropy method, uniform Sobolev estimates and compactness arguments, which allow us to rigorously justify the convergence of solutions of the compressible system toward the limit system as $\varepsilon\to 0$. Our work provides a mathematical foundation for the incompressible regime of two-phase models with algebraic closure and single velocity, and clarifies the structure of the low Mach number limit in this context.
This paper is organized as follows. In Section 2, we introduce our relative entropy functional which consists of a classical contribution and an additional component based on a logarithmic comparison of the partial densities. In Section 3, we exploit the special structure of the bi-fluid system to derive several key inequalities, including an original comparison between the $L^{1}$ and $L^{2}$ norms of the partial densities and a variant of the Csiszár–Kullback–Pinsker inequality. Next, in Section 4, we prove the strong convergence of the partial densities, and finally, Section 5 completes the proof of our main theorem.
2 Relative entropy
In this section, we introduce the relative entropy functional that will be our main tool to prove Theorem 1.6.
2.1 Definition
We derive a relative entropy between a weak solution $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ of System (LABEL:system_eps), in the sense of Definition 2.1 below, and any other state $(\rho_{+},\rho_{-},\alpha_{+},u)$ of the fluid. Let us consider the relative energy functional, denoted by ${\mathcal{E}}_{1}(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u)$ and defined as:
$\begin{split}{\mathcal{E}}_{1}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)=&\int_{\Omega}\frac{1}{2}(\alpha^{\varepsilon}_{+}\rho^{\varepsilon}_{+}+\alpha^{\varepsilon}_{-}\rho^{\varepsilon}_{-})|u^{\varepsilon}-u|^{2}\\ &+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}}H(\rho_{+}^{\varepsilon}|\rho_{+})+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}}H(\rho_{-}^{\varepsilon}|\rho_{-}),\end{split}$ (2.1)
with, for $i=+,$-,
$H(\rho_{i}^{\varepsilon}|\rho_{i}):=\frac{1}{\gamma_{i}-1}\left((\rho_{i}^{\varepsilon})^{\gamma_{i}}-\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\right).$ (2.2)
In order to control the $L^{1}$ norm of the partial densities and close the Gronwall argument required for the proof of Theorem 1.6, we introduce the following functional
[…]
and we set $\ln_{+}s=\max(0,\ln s)$.
Finally, we define the total relative entropy functional:
${\mathcal{E}}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)=({\mathcal{E}}_{1}+{\mathcal{E}}_{2})\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big).$
In this definition, we always set $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}$ and $\alpha_{-}=1-\alpha_{+}$.
The goal of the next subsection is to study the evolution of this entropy for the weak solution of System (LABEL:system_eps) in the following sense.
@@ 2.1 Definition
Let us consider some initial data $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ for $\gamma_{\pm}\geqslant 2$ satisfying the compatibility conditions (1.4), where we defined $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$ from $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})$ (see Remark 1.2). Let $T>0$. We say that $(\rho_{\pm}^{\varepsilon},\alpha_{\pm}^{\varepsilon},u^{\varepsilon})$ is a finite energy weak solution of the compressible two-phase Navier-Stokes system (LABEL:system_eps) with algebraic closure on $(0,T)$ associated with the initial data $(\rho_{\pm,0}^{\varepsilon},\alpha_{\pm,0}^{\varepsilon},u_{0}^{\varepsilon})$ if we have
@@
    $0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\quad\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\quad\rho_{\pm}^{\varepsilon}\geqslant 0,\quad\rho_{+}^{\gamma_{+}}=\rho_{-}^{\gamma_{-}}\text{ a.e. in }[0,T]\times\Omega,$
@@
    $\int_{\Omega}(\rho^{\varepsilon}u^{\varepsilon})(t,x)\cdot\phi(t,x)dx$ $-\int_{\Omega}(\rho^{\varepsilon}_{0}u^{\varepsilon}_{0}(x)\cdot\phi(0,x)dx$
    $=$ $\int_{0}^{t}\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot\partial_{t}\phi+\rho^{\varepsilon}(u^{\varepsilon}\otimes u^{\varepsilon}):\nabla\phi+\frac{1}{\varepsilon^{2}}(\rho_{+}^{\varepsilon})^{\gamma_{+}}\operatorname{div}\phi\Big)(s,x)dxds$
    $-\int_{0}^{t}\int_{\Omega}\left(\mu\nabla u^{\varepsilon}:\nabla\phi+(\mu+\lambda)\operatorname{div}u^{\varepsilon}\operatorname{div}\phi\right)(s,x)dxds$
@@
    $E(t)+\mu\int_{0}^{t}\int_{\Omega}\left(\nabla u^{\varepsilon}\right)^{2}\,\mathrm{d}x\,\mathrm{d}s+\left(\lambda+\mu\right)\int_{0}^{t}\int_{\Omega}(\operatorname{div}u^{\varepsilon})^{2}\,\mathrm{d}x\,\mathrm{d}s\leqslant E(0)$ (2.4)
    for a.e. $t\in(0,T)$, where
    $E(t)=\int_{\Omega}\frac{1}{2}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(t,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon\gamma_{+}}(t,x)}{\varepsilon^{2}(\gamma_{+}-1)}\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon\gamma_{-}}(t,x)}{\varepsilon^{2}(\gamma_{-}-1)}\,\mathrm{d}x$ (2.5)
@@
The functional appearing in the energy inequality is the natural energy associated with the system. A detailed formal derivation of the corresponding energy identity for smooth solutions can be found in [29].

### v3

@@ 1.1 Main result
It should be possible to extend this analysis to periodic domains or to two-dimensional settings.
1.2 Survey of previous results
Compressible two-fluid models have been extensively studied in the literature. From a mathematical perspective, we refer for instance to the survey chapter by Bresch, Desjardins, Ghidaglia, Grenier and Hillairet [6] and the references therein. We briefly review below some results that are most relevant to the present work.
In the viscous setting, the analysis of such systems has led to important developments in the theory of compressible PDEs. In particular, the existence of finite-energy weak solutions was established by Bresch, Mucha, and Zatorska [10] as well as by Vasseur, Wen, and Yu [57], highlighting suitable dissipation mechanisms. In the same spirit, Novotný and Pokorný [49] and Novotný [48] developed a weak solution theory of compressible bi-fluid models, while Jin, Kwon, Nečasová, and Novotný [31] introduced dissipative turbulent solutions and investigated their stability. Weak-strong uniqueness properties were further analyzed in [32, 40, 43]. These contributions, among others, as e.g. [39, 42, 51], form a coherent and evolving theory for compressible two-fluid flows.
For single-phase flows, the low Mach number limit is by now well understood. For the compressible Euler and Navier–Stokes equations, convergence toward incompressible dynamics under well-prepared initial data was first established by Klainerman and Majda [34, 35]. This analysis was later extended by Schochet [53], who developed a general framework to handle fast oscillations in hyperbolic systems.
In the isentropic setting, numerous results are available. In the framework of weak solutions, the low Mach number limit for the compressible Navier–Stokes system was established in [46], [20], [19], and [7]. For strong solutions, we refer to [28] and [15].
In the non-isentropic case, the analysis is more involved due to the coupling with the temperature (or energy) equation. The low Mach number limit has been justified in the weak solution framework in [24] and [26] (see also the books [45] and [22]). For strong solutions, results were obtained for instance in [47], [2], and [1].
A concise overview of these developments can be found in [27].
Let us also mention that the incompressible limit for the classical isentropic compressible Navier–Stokes system may be obtained in other asymptotic regimes. In particular, the large viscosity limit leading to the incompressible equations was studied by Danchin and Mucha [18]. We may also refer to Fanelli et al [21] for related incompressible limits in a reduced compressible MHD system.
A comprehensive treatment of the compressible Navier–Stokes system with constant viscosities, with particular emphasis on the relative entropy method as a stability and convergence tool, is given in [25]. For more complex systems, such as compressible Navier–Stokes equations with density-dependent viscosities, the notion of relative entropy had to be revisited and adapted; see, for instance, [11] and [8]. We also mention works developing the relative entropy framework in related contexts, such as [23] and [55].
In contrast, significantly fewer results are available for multiphase flow models, even in simplified configurations. Only recently has attention been devoted to the formal derivation of low Mach number limits for various two-phase systems, including models with algebraic or PDE closures, as well as formulations involving one or two velocities. In these works, asymptotic expansions are typically performed as the Mach number tends to zero; see, for instance, [56] and [37]. Numerical investigations of such regimes have also been carried out, see [4] and the references therein.
In the same week that we posted the first version of this paper on ArXiv, an independent work [40] also appeared, dealing with the bi-fluid compressible Navier–Stokes equations. In that paper, the authors investigate weak–strong uniqueness through the introduction of a different relative entropy. Although their functional contains the same natural term $\mathcal{E}_{1}^{\varepsilon}$, the additional contribution is simply the $L^{2}$-distance between the volume fractions $\alpha_{i}^{\varepsilon}-\alpha_{i}$. This choice leads to an entropy of a completely different nature from our functional $\mathcal{E}_{2}^{\varepsilon}$ defined in (1.8), which is built solely on the partial densities $R_{i}^{\varepsilon}$. Let us emphasize that the partial densities are the natural variables for the bi-fluid compressible Navier–Stokes system with algebraic closure (LABEL:system_eps), as highlighted in the majority of the literature (see, among others, [10, 49]).
The month after our initial submission, another related paper appeared on ArXiv: in [41], the authors reused the entropy introduced in [40] to obtain a low-Mach number limit under very well-prepared initial data. Besides assuming that $\rho_{i}^{\varepsilon}(0)$ is close to the constant $C_{0}^{1/\gamma_{i}}$—a condition encoded by the uniform bound of $\varepsilon^{-2}H_{i}(\rho_{i,0}^{\varepsilon})$, which is also present in our assumptions and is standard in low-Mach analyses—they further require that $\alpha_{i}^{\varepsilon}(0)$ is close to the constant $\overline{\alpha_{i}}=1/C_{0}^{\gamma_{i}}$ in high Sobolev norms. The additional constraint $\overline{\alpha_{+}}+\overline{\alpha_{-}}=1$ imposes a strong restriction on the admissible choices of $C_{0}$. In this setting, the authors naturally recover in the limit the homogeneous incompressible Navier–Stokes equations, characterized by the constant density $\rho=\overline{\alpha_{+}}\rho_{+}+\overline{\alpha_{-}}\rho_{-}=2$. Although studying this particular regime—where the volume fractions remain close to constants—is an interesting and meaningful first step (and is included in our approach), the primary purpose of our entropy functionals is precisely to address genuinely non-trivial mixtures. Our framework handles the full complexity of the bi-fluid model, in which the coupling between $\alpha_{i}$ and $R_{i}$ is highly non-linear and non-trivial.
For completeness, we should also emphasize that the analysis in [41] is entirely independent of ours. Instead of working with global weak solutions of the compressible system, the authors consider local strong solutions on a time interval $T_{\varepsilon}$ that is smaller than $T$, the existence time of the solution to the limit system. Their analysis should be interpreted as a perturbative analysis around constant states $(\rho_{i},\overline{\alpha_{i}})$, allowing in particular $\gamma_{\pm}>1$.
1.3 Strategy of the proof
The proof is based on the relative entropy method. More precisely, we consider the modulated quantities introduced above, which measure the deviation between a solution of the compressible system and a strong solution of the limit system, and combine them into a relative entropy functional. The main argument then consists in deriving a stability inequality for this functional and applying a Gronwall-type argument to conclude the convergence as $\varepsilon\to 0$.
The strategy is inspired by the approach of Lions and Masmoudi [46]. However, the two-phase structure of the system introduces additional difficulties. In particular, the standard relative entropy does not provide sufficient control to close the argument, as it fails to yield bounds on the $L^{2}$-distance between $R_{\pm}^{\varepsilon}$ and $R_{\pm}$. To overcome this issue, we introduce a logarithmic relative entropy, which provides the required control on the densities and their deviations from the limit state. Combined with the modulated entropy, which encodes the natural stability structure of the system, this yields a suitable framework for the analysis.
With these tools in hand, we proceed in three main steps.
Step 1: Uniform bounds via a first Gronwall argument. We first establish uniform bounds for ${\mathcal{E}}_{1}^{\varepsilon}$ and ${\mathcal{E}}_{2}^{\varepsilon}$. The key observation is that, while each functional alone does not yield sufficient control, their sum does. By deriving a differential inequality for ${\mathcal{E}}_{1}^{\varepsilon}+{\mathcal{E}}_{2}^{\varepsilon}$ and applying a first Gronwall argument, we obtain uniform bounds on both ${\mathcal{E}}_{1}^{\varepsilon}$ and ${\mathcal{E}}_{2}^{\varepsilon}$ individually.
Step 2: Strong convergence of densities and weak convergence of the divergence. The uniform bound on ${\mathcal{E}}_{1}^{\varepsilon}$ yields the strong convergence of the densities $\rho_{\pm}^{\varepsilon}$. In addition, combining the control of ${\mathcal{E}}_{1}^{\varepsilon}+{\mathcal{E}}_{2}^{\varepsilon}$ with compactness arguments allows us to deduce the weak convergence of $\operatorname{div}u^{\varepsilon}$ to 0.
Step 3: Strong convergence of the velocity and conclusion. Finally, a second Gronwall argument ensures the strong convergence of the velocity field $u^{\varepsilon}$ as well as that of the partial densities. Next, the convergence of the volume fractions relies crucially on the specific structure of the bi-fluid system.
This closes the argument, as the convergence of all relevant quantities is then established, and the limit system is identified.
This paper is organized as follows. Section 2 is devoted to the mathematical setting of the problem and recalls several preliminary results. In Section 3, we establish some inequalities for the relative entropies. Next, in Section 4, we find some useful uniform bounds for the relative entropies using a first Gronwall estimate. This allows us to establish the strong convergence of the densities and the weak divergence of $\operatorname{div}u^{\varepsilon}$. Finally, Section 5 completes the proof of our main theorem by proving the strong convergences of the velocity field and the volume fractions. Appendix A provides entropy computations for more general velocity fields $u$, which may be useful for analyzing stability or establishing weak–strong uniqueness.
2 Preliminaries and mathematical setting
2.1 Bi-fluid compressible system and nondimensionalization
In the present paper, we analyse a bi-fluid isentropic compressible Navier-Stokes system for barotropic pressures. In this subsection, we present this system and its nondimensionalization. Denoting the phases by ++ and -, we are considering the following system.
where $\rho_{i}$, $p_{i}$, $\alpha_{i}$ the density, pressure and volume fraction of phase $i$. Let us notice that, thanks to the algebraic closure, the pressure term in the momentum equation can be simplified:
The system is endowed with the Dirichlet boundary condition (1.2) for $u(t,x)$ on $(0,T)\times\partial\Omega$.
The incompressible limit is the limit where the speed of the fluid becomes negligible compared to the speed of sound. Thus we reset the flow speed to the scale by replacing $u$ with $\varepsilon u^{\varepsilon}$. A particle will move a distance of order $1/\varepsilon$ during a time of order $1/\varepsilon^{2}$, meaning that the relevant time scale is $1/\varepsilon^{2}$. This is why we can use a change of variable like $t^{\varepsilon}=\varepsilon^{2}t$ and $x^{\varepsilon}=\varepsilon x$, with $t^{\varepsilon}$ and $x^{\varepsilon}$ of order $1$ (see [16]). Thus, we scale $u$ and $\rho$ in the following way
Observe that, under this change of variables, the viscosity coefficients do not require rescaling.
Therefore, System (LABEL:system) can be rewritten as System (LABEL:system_eps).
2.2 Definition of a weak solution of the compressible bi-fluid system
We consider weak solutions to System (LABEL:system_eps) in the following sense.
In this definition, we always set $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}$.
@@ 2.2 Definition of a weak solution of the compressible bi-fluid system
Let us consider some initial data $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}(\Omega)\times L^{\gamma_{+}}(\Omega)\times L^{1}(\Omega)$ for $\gamma_{\pm}\geqslant 2$ satisfying the compatibility conditions (1.3), where we defined $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$ from $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})$ (see Remark 1.1). Let $T>0$. We say that $(\rho_{\pm}^{\varepsilon},\alpha_{\pm}^{\varepsilon},u^{\varepsilon})$ is a finite energy weak solution of the compressible two-phase Navier-Stokes System (LABEL:system_eps) with algebraic closure on $(0,T)$ associated with the initial data $(\rho_{\pm,0}^{\varepsilon},\alpha_{\pm,0}^{\varepsilon},u_{0}^{\varepsilon})$ if we have
@@
    $0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\quad\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\quad\rho_{\pm}^{\varepsilon}\geqslant 0,\quad\rho_{+}^{\gamma_{+}}=\rho_{-}^{\gamma_{-}}\text{ a.e. in }(0,T)\times\Omega,$
@@
    $\int_{\Omega}(\rho^{\varepsilon}u^{\varepsilon})(t,x)\cdot\phi(t,x)\,\mathrm{d}x$ $-\int_{\Omega}(\rho^{\varepsilon}_{0}u^{\varepsilon}_{0}(x)\cdot\phi(0,x)\,\mathrm{d}x$
    $=$ $\int_{0}^{t}\!\!\!\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot\partial_{t}\phi+\rho^{\varepsilon}(u^{\varepsilon}\otimes u^{\varepsilon}):\nabla\phi+\frac{1}{\varepsilon^{2}}(\rho_{+}^{\varepsilon})^{\gamma_{+}}\operatorname{div}\phi\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
    $-\int_{0}^{t}\!\!\!\int_{\Omega}\left(\mu\nabla u^{\varepsilon}:\nabla\phi+(\mu+\lambda)\operatorname{div}u^{\varepsilon}\operatorname{div}\phi\right)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
@@
    $E^{\varepsilon}(t)+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla u^{\varepsilon}\right)^{2}\,\mathrm{d}x\,\mathrm{d}s+\left(\lambda+\mu\right)\int_{0}^{t}\!\!\!\int_{\Omega}(\operatorname{div}u^{\varepsilon})^{2}\,\mathrm{d}x\,\mathrm{d}s\leqslant E^{\varepsilon}(0)$ (2.2)
    for a.e. $t\in(0,T)$.
@@
The above inequality follows from the classical energy inequality (see, e.g., [49]) and the special structure of the functional $H$, see the next section.
2.3 Energy conservation
In this section, we provide some details of the formal computation leading to the energy conservation for completeness. We first refer to [37] for the terms whose contribution to the energy balance has already been derived
$\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(t,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}(t,x)}{\varepsilon^{2}}\frac{\rho_{+}^{\varepsilon\gamma_{+}}}{\gamma_{+}-1}(t,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}(t,x)}{\varepsilon^{2}}\frac{\rho_{-}^{\varepsilon\gamma_{-}}}{\gamma_{-}-1}(t,x)\,\mathrm{d}x$
$\quad\int_{0}^{t}\int_{\Omega}\mu|\nabla u|^{2}(t,x)+(\mu+\lambda)|\operatorname{div}u|^{2}(t,x)\,\mathrm{d}x$
$\quad\leqslant\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(0,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}(0,x)}{\varepsilon^{2}}\frac{\rho_{+}^{\varepsilon\gamma_{+}}}{\gamma_{+}-1}(0,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}(0,x)}{\varepsilon^{2}}\frac{\rho_{-}^{\varepsilon\gamma_{-}}}{\gamma_{-}-1}(0,x)\,\mathrm{d}x.$
We now focus on the remaining terms and show that they do not contribute to the energy variation:
$-\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\Big(\rho_{+}^{\gamma_{+}}+\gamma_{+}\rho_{+}^{\gamma_{+}-1}(\rho_{+}^{\varepsilon}-\rho_{+})\Big)\,\mathrm{d}x-\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\Big(\rho_{-}^{\gamma_{-}}+\gamma_{-}\rho_{-}^{\gamma_{-}-1}(\rho_{-}^{\varepsilon}-\rho_{-})\Big)\,\mathrm{d}x$
$=\frac{\rho_{+}^{\gamma_{+}}}{\varepsilon^{2}}\int_{\Omega}\alpha_{+}^{\varepsilon}\,\mathrm{d}x-\frac{\gamma_{+}\rho_{+}^{\gamma_{+}-1}}{\varepsilon^{2}(\gamma_{+}-1)}\int_{\Omega}\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}\,\mathrm{d}x+\frac{\rho_{-}^{\gamma_{-}}}{\varepsilon^{2}}\int_{\Omega}\alpha_{-}^{\varepsilon}\,\mathrm{d}x-\frac{\gamma_{-}\rho_{-}^{\gamma_{-}-1}}{\varepsilon^{2}(\gamma_{-}-1)}\int_{\Omega}\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\,\mathrm{d}x$
$=\frac{\rho_{+}^{\gamma_{+}}}{\varepsilon^{2}}\int_{\Omega}1\,\mathrm{d}x-\frac{\gamma_{+}\rho_{+}^{\gamma_{+}-1}}{\varepsilon^{2}(\gamma_{+}-1)}\int_{\Omega}R_{+}^{\varepsilon}\,\mathrm{d}x-\frac{\gamma_{-}\rho_{-}^{\gamma_{-}-1}}{\varepsilon^{2}(\gamma_{-}-1)}\int_{\Omega}R_{-}^{\varepsilon}\,\mathrm{d}x,$
where we use the equality of pressures $\rho_{+}^{\gamma_{+}}=\rho_{-}^{\gamma_{-}}=C_{0}$ and $\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1$. We can conclude using the conservation of total mass due to the mass equation in System (LABEL:system_eps):
$\int_{\Omega}R_{\pm}^{\varepsilon}(t,x)\,\,\mathrm{d}x=\int_{\Omega}R_{\pm,0}^{\varepsilon}(x)\,\,\mathrm{d}x\qquad\text{for all }t\in(0,T).$ (2.3)
2.4 Non-homogeneous incompressible system
In the case of nonhomogeneous incompressible flows with constant viscosities, the existence of both weak and strong solutions has been investigated in several works. The existence of weak solutions was first established by Kazhikhov [33]. Concerning strong solutions, Ladyzhenskaya and Solonnikov [36] proved local existence, as well as global existence for small initial data (and global solvability in two space dimensions). In sufficiently smooth domains, Lemoine [38] obtained local strong regular solutions, which are global for small regular data. Okamoto [50] proved existence and uniqueness of local weak solutions and obtained global solutions in three dimensions under smallness assumptions on the initial data. Simon [54] established the existence of global-in-time weak solutions under suitable weak assumptions on the initial data, and proved global strong solvability provided the initial density is not too small (and sufficiently regular). The solvability of nonstationary problems for nonhomogeneous incompressible fluids, as well as convergence in the vanishing viscosity limit, was studied by Itoh [30]. Choe and Kim [14] established local existence and uniqueness of strong solutions in three dimensions, including situations where the initial density vanishes on an open subset of the domain. More recently, Danchin [17] developed a well-posedness theory in critical Besov spaces and proved the existence of global-in-time unique solutions under suitable smallness assumptions on the initial data.
Among the vast existing literature, let us state a result of local existence of strong solutions of non-homogeneous viscous incompressible Navier-Stokes equations, coming from [52].
Theorem 2.3.
Let $V=\{u\in H^{1}(\Omega)\text{ such that }\operatorname{div}u=0\text{ and }u\cdot n|_{\partial\Omega}=0\}$. Let the boundary $\partial\Omega$ be of class $C^{\infty}$. Let $(\rho_{0},u_{0})$ be some initial data such that $u_{0}\in H^{3}(\Omega)\cap V$ in $\Omega$ and $\rho_{0}\in C^{\infty}(\bar{\Omega})$, with $\overline{C}\geqslant\rho_{0}\geqslant\underline{C}>0$ for some constants $\overline{C},\underline{C}$. Then there exists $T>0$ and a unique solution $(\rho,u,\Pi)$ of System (1.4) that satisfy
$u\in L^{\infty}(0,T;V)\cap C(0,T;H^{3}(\Omega)\cap V),\qquad\partial_{t}u\in C(0,T;H^{1}(\Omega)\cap V)\cap L^{2}([0,T]\times\Omega),$
$\rho\in C^{1}([0,T]\times\Omega),\qquad\partial_{t}\rho\in L^{\infty}(0,T;H^{-1}(\Omega)),$
$\nabla\Pi\in L^{\infty}(0,T;L^{3}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega)).$
Remark 2.4.
which is well defined for $\gamma_{+}\neq\gamma_{-}$, as $C_{0}>0$. We note that $\alpha_{\pm}$ satisfy a transport equation. If the initial data verify $\alpha_{\pm,0}\in[\underline{a},\overline{a}]\subset(0,1)$, then this property is preserved for all times. Likewise, since $\rho$ also satisfies a transport equation, it follows that $0<\underline{C}\leqslant\rho\leqslant\overline{C}$. Moreover, the constants $\underline{C}$ and $\overline{C}$ can be expressed in terms of $\underline{a}$, $\overline{a}$, $C_{0}$, and $\gamma_{\pm}$.
Remark 2.5.
The only difference compared to Theorem 3 of Salvi is that we additionally assume that $\nabla\Pi$ has the same regularity as $\Delta u$. We do not claim that this regularity is optimal for our analysis; however, we rely at some point on the fact that $\Delta u$ and $\nabla\Pi$ belong to $L^{\infty}(0,T;L^{3}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$, as guaranteed by Theorem 2.3.
3 Inequalities for the relative entropies
In this section, we establish some inequalities on the relative entropies defined in (1.5)-(1.8) which will allow us to prove Theorem 1.3.
In the following, $C$ denotes a generic constant which may depend, depending on the context, on $\underline{C},\overline{C},\underline{a},\overline{a},C_{0}$ or $\gamma_{\pm}$.